# Non-local Metric-Affine Gravity

### v1

@@ I Introduction
It is widely believed that quantum gravitational effects manifest through non-locality. This feature is invoked in order to reconcile General Relativity (GR) with Quantum Field Theory (QFT) prescriptions. These quantum corrections, that alter the classical gravitational action, encode non-locality and dramatically change the associated dynamics producing new degrees of freedom. Non-locality is typically expressed through the inclusion of inverse powers of the box operator $\Box^{-1}$ sandwiched between geometric quantities (like the Ricci scalar) that supplement the gravitational action and consequently the associated field equations Mashhoon:2017qyw ; Deser:2007jk ; Capozziello:2008gu ; Nojiri:2010pw . Another kind of non-locality is realized by adding an infinite series of the box operator in the form of entire functions like e.g. $RF(\Box)R$ where $R$ is the Ricci scalar and $F(\Box)$ a holomorphic function of the box operator. This is known as Infinite Derivative Gravity (IDG) efimov1967non . Besides fundamental physics issues, the key idea to adopt non-local gravity models is to obtain characteristic lengths and masses capable of addressing Ultra-Violet and Infra-Red issues in astrophysics and cosmology Blome:2010xn ; Roshan:2022zov .
@@
Cosmological applications of non-local gravity have been studied in Refs. Deser:2007jk ; Koivisto:2008xfa ; Calcagni:2018pro ; Capozziello:2024qol ; Capozziello:2021krv ; Nojiri:2019dio ; Biswas:2010zk ; Kolar:2021qox ; Koshelev:2016xqb ; Calcagni:2013vra ; Dirian:2014bma and some problems like the Weinberg no-go theorem for cosmological constant can be addressed in this framework Capozziello:2025bsm . The teleparallel formulation of non-local gravity along with its cosmological applications, has been considered in Bahamonde:2017bps . Also black hole solutions can be improved with the presence of non-local terms Bambi:2016uda ; Capozziello:2025iwn as well as gravitational waves where new polarizations emerge Capozziello:2024bjb ; Foffa:2013sma .
@@
Here we are going to formulate non-local gravity in the metric-affine geometry setup in order to consider the most general framework where these theories can be formulated. This approach proves to be useful also in view of extensions and modifications of General Relativity where the geometric invariants can be other than the standard curvature Bahamonde:2017bps ; BeltranJimenez:2019esp ; Capozziello:2022zzh ; Capozziello:2025hyw ; Golovnev:2023yla .
@@
The paper is organized as follows. In Sec.IIII, we introduce the notation and conventions we are going to use. Then, in Sec. $III$, we start with the metric-affine analogue of the theory introduced in Ref. Gottlober:1989ww , which was studied in the Riemannian setting. Immediately the differences with the metric formalism emerge. By solving the connection field equations of the theory we cast it in an equivalent Scalar-Tensor metric theory with higher order derivative terms for the scalar field. Consequently we generalize it in Sec. $IV$ and study the metric-affine version of the general class of theories with the Lagrangians $\mathcal{L}=RF(\Box)R$ and $\mathcal{L}=RF(\Box^{-1})R$. Deriving the field equations, and solving the connection equations, we again establish equivalence with a generalized higher-derivative Scalar-Tensor theory over a Riemannian background. Finally, we introduce a new non-local Metric-Affine Gravity theory which has no Riemannian analogue and consists of the newly introduced terms $X\Box X$, where $X$ stands for torsion and non-metricity, into the action, along with the usual Einstein-Hilbert term. Obtaining the field equations we show how torsion and non-metricity are dynamical fields in this setting even in vacuum. Some applications to cosmology are discussed in Sec $VI$. In Sec. $VII$, we then wrap up our results and point to future directions.
@@ II The Metric-Affine formalism
The difference between the general connection $\Gamma^{\lambda}{}_{\mu\nu}$ and the Levi-Civita one, $\tilde{\Gamma}^{\lambda}{}_{\mu\nu}$, defines a tensor that quantifies how much the geometry deviates from being Riemannian. This is the so-called distortion tensor schouten1954ricci ; hehl1995metric
@@
This is called the post-Riemannian expansion of the connection and it is particularly useful for splitting the various quantities into their Riemannian parts [^1] Namely the ones computed with respect to the Levi-Civita connection. plus the non-Riemannian contributions of torsion and non-metricity. Torsion and non-metricity can then be recovered from the distortion from the relations (see e.g. Hehl:1994ue ),
@@
In this form, it can further act upon tensor fields of arbitrary rank, $\Box T^{\mu_{1}...\mu_{p}}{}{}_{\nu_{1}...\nu_{q}}=\widetilde{\nabla}^{\mu}\widetilde{\nabla}_{\mu}T^{\mu_{1}...\mu_{p}}{}{}_{\nu_{1}...\nu_{q}}$. Note that in Briscese:2015urz the definition
@@
where $\nabla$ is the full covariant derivative associated with the full connection $\Gamma^{\lambda}{}_{\mu\nu}$, was used in a Palatini formulation of non-local gravity. Although this could be a viable option, there are several reasons to avoid such a definition in the current context. First and foremost, (10) does not reduce to (8) when acted upon functions. Secondly, given the fact that the metric is not covariantly constant with respect to the full connection, the above definition is not the only one that reduces to $\partial^{\mu}\partial_{\mu}$ in the flat space limit. Indeed, operators like $\nabla_{\mu}(g^{\mu\nu}\nabla_{\nu})$, $\frac{1}{\sqrt{-g}}\nabla_{\mu}(\sqrt{-g}g^{\mu\nu}\nabla_{\nu})$ or $\nabla_{\mu}(\nabla_{\nu}g^{\mu\nu})$ also reduce to $\partial^{\mu}\partial_{\nu}$ in the flat space limit. One would then need to take all possible linear combinations of such operators (with the coefficients added up to one) in order not to discriminate among them, see for instance Shimada:2018lnm . However, this approach, not only does it complicate the analysis but also introduces new arbitrary parameters. Thirdly, for operators like (10), terms like $R_{\mu\nu}\Box^{(\Gamma)}R^{\mu\nu}$ and $R^{\mu\nu}\Box^{(\Gamma)}R_{\mu\nu}$ do not coincide since the metric cannot freely move through the full covariant derivative $\nabla_{\mu}$. This then complicates the analysis even further since one would be forced to include both of them in order to maintain full generality. Finally, the basic property of the inner product $(\phi\Box\psi)=(\Box\phi,\psi)$ is lost if one uses (10) instead of (8). It is therefore only natural to stick to (8) for the definition of the box operator in our current context.
@@ III A simple Warm-up Theory
Let us start with the model in Gottlober:1989ww , but now in the metric-affine formalism, where the connection and the metric are independent fields. We have
@@
where $\alpha$ is dimensionless [^2] Of course, the General Relativity limit is recovered for $\alpha=1$. and $\beta$ and $\gamma$ are parameters with mass dimension 2-2 and 4-4 respectively. Also, we define the box operator associated to the metric $g$, in the usual way, (see eq. (8)) as we explained in the previous section. Let us also note that, in this action, the scalar curvature $R$ is now the generic one that has also torsion and non-metricity. The action (11) is, up to surface terms, equivalent to
@@
Using the method developed in Iosifidis:2018jwu , from the above connection field equations, one finds the distortion
@@
Now notice that our action (11) is invariant under projective transformations of the connection
@@
since it is built solely from $R$ which is itself projectively invariant. This means that there is an undetermined vectorial degree of freedom which can be fixed by a gauge choice iosifidis2019exactly . Two special gauge choices are of interest here. The first one is to fix
@@
This is the vanishing torsion gauge. We see therefore a torsion/non-metricity duality similar to the one that appears in $f(R)$ theories Iosifidis:2018zjj . Let us now investigate the metric field equations (13). Taking the trace of the latter, it follows that
@@
with $\phi=1+2\gamma\Box\psi$. We see therefore that the initial theory is on-shell equivalent to a Scalar-Tensor theory with higher order derivatives for the scalar field. It is important to stress that, in constrast to the Riemannian case where the resulting metric field equations contain sixth order derivatives of the scalar field, in the metric-affine approach developed here, we have up to fourth-order derivatives of the scalar field. In order words, with respect to the metric formalism developed in Gottlober:1989ww , the metric-affine approach reduces the dynamics. See also Amendola:1993bg for a cosmological application.
@@ IV.1 Nonlocal Ultraviolet $\mathcal{L}=R+RF(\Box)R$ Theories
where $F(\Box)$ is an analytic function of the box operator, usually called the form factor BasiBeneito:2022wux , which we may expand as
@@
In the first representation $M^{2}$ carries a squared mass scale and consequently the coefficients $a_{m}$ are dimensionless constants. In the second form, the mass dependence is absorbed in $f_{m}$. In order to have more simplified expressions, we shall use the latter representation. Non-local theories of this kind, namely consisting of positive powers of the box operator are relevant in the UV regime Buoninfante:2022trw . In the metric formalism, it was shown in Biswas:2005qr that the class (48) gives rise to ghost and asymptotically free theories. Below we shall analyze the dynamics of these theories in the metric-affine approach.
@@
Note that these are formally the same with (14), with the mere replacement $\gamma\Box R\rightarrow F(\Box)R$ and of course setting $\alpha=1$ and $\beta=0$. It is therefore again possible to solve for the connection. Indeed, despite the complexity of the metric field equations, the connection field equations obtained above, have a rather simple structure. As a matter of fact, since the form of (45) is identical to that of (14) with only the exact form of $\phi$ altered, the exact and unique solution to (45) is still
@@ IV.2 Non-local Infrared $\mathcal{L}=R+RF(\Box^{-1})R$ Theories
The previous results can be straightforwardly applied to non-local theories involving the inverse box operator Deser:2007jk $\Box^{-1}$ defined through $\Box\Box^{-1}=\Box^{-1}\Box=\mathbb{1}$. As a prototype let us consider the class of non-local theories given by [^3] For the study of such theories in the metric formalism see e.g. Capozziello:2020xem ; Capozziello:2021bki ; Capozziello:2023mju .
@@
where $\Box^{-m}$ is an abbreviation for $(\Box^{-1})^{m}$. Note that this form of series for $F$, namely containing powers of the inverse box operator are relevant for the IR regime. In order to proceed with the field equations, we only need to vary the defining relation of $\Box^{-1}$ and subsequently multiply from the left by the same operator, to trivially get
@@ V Novel Theories with no Riemannian analogue
As we have already pointed out, the action (11) as well as its extension (48), have been well studied in the Riemannian case. In particular, in a Riemannian setting, (11) is the most general theory that is quadratic in scalar curvature. We note, however, that in MAG, squared torsion and non-metricity terms supplement the linear in curvature Einstein-Hilbert term. Therefore, in a MAG scheme before extending to (11) it would be quite natural to consider non-local actions consisting of scalars formed by sandwiching the box operator with torsion and non-metricity before moving to quadratic curvature. Given that there are 11 parity even quadratic invariants, the analogue of the theory (11) to MAG is
@@
Some comments are in order. The first two lines consist of the usual Einstein-Hilbert term along with the quadratic invariants in torsion and non-metricity. These exhaust the number of parity even invariants up to dimension $L^{-2}]$. When only the first two lines of the above action are considered then the vacuum theory reduces to General Relativity Iosifidis:2021tvx . The third and fourth lines are the new non-local terms we consider. These have not been studied previously in the literature. Let qualitatively study their dynamics. The full field equations are quite lengthy and are given in the appendix. Suppressing indices, the connection field equations schematically look like
@@
where $N$ is the distortion. This shows that torsion and non-metricity propagate in this case and the theory has more degrees of freedom than General Relativity. It goes beyond the scope of the current study to investigate the healthy sub-sectors of this class of theories. However, it is worth mentioning that the non-local terms of the third and fourth lines, when added to the quadratic curvature MAG theories, may help to extend the healthy spectrum of the quadratic MAG. In order to get our point across, let us study a simple, but yet non-trivial, subsector of the above class.

### v2

@@ I Introduction
It is widely believed that quantum gravitational effects manifest through non-locality. This feature is invoked in order to reconcile General Relativity (GR) with Quantum Field Theory (QFT) prescriptions. These quantum corrections, that alter the classical gravitational action, encode non-locality and dramatically change the associated dynamics producing new degrees of freedom. Non-locality is typically expressed through the inclusion of inverse powers of the box operator $\Box^{-1}$ sandwiched between geometric quantities (like the Ricci scalar) that supplement the gravitational action and consequently the associated field equations Mashhoon (2017); Deser and Woodard (2007); Capozziello et al. (2009); Nojiri et al. (2011). Another kind of non-locality is realized by adding an infinite series of the box operator in the form of entire functions like e.g. $RF(\Box)R$ where $R$ is the Ricci scalar and $F(\Box)$ a holomorphic function of the box operator. This is known as Infinite Derivative Gravity (IDG) Efimov (1967). Besides fundamental physics issues, the key idea to adopt non-local gravity models is to obtain characteristic lengths and masses capable of addressing Ultra-Violet and Infra-Red issues in astrophysics and cosmology Blome et al. (2010); Roshan and Mashhoon (2022); Boos (2020); Boos et al. (2018).
@@
Cosmological applications of non-local gravity have been studied in Refs. Deser and Woodard (2007); Koivisto (2008); Calcagni et al. (2018); Capozziello et al. (2024); Capozziello and Bajardi (2022); Nojiri et al. (2020); Biswas et al. (2010); Kolář et al. (2022); Koshelev et al. (2016); Calcagni et al. (2014); Dirian et al. (2015) and some problems like the Weinberg no-go theorem for cosmological constant can be addressed in this framework Capozziello et al. (2025c). Non-local hybrid metric-Palatini cosmology was studied in Bombacigno et al. (2025). The teleparallel formulation of non-local gravity along with its cosmological applications, has been considered in Bahamonde et al. (2017). Also black hole solutions can be improved with the presence of non-local terms Bambi et al. (2016); Capozziello et al. (2025d) as well as gravitational waves where new polarizations emerge Capozziello et al. (2025a); Foffa et al. (2014).
@@
Here we are going to formulate non-local gravity in the metric-affine geometry setup in order to consider the most general framework where these theories can be formulated. This approach proves to be useful also in view of extensions and modifications of General Relativity where the geometric invariants can be other than the standard curvature Bahamonde et al. (2017); Beltrán Jiménez et al. (2019); Capozziello et al. (2022, 2025b); Golovnev (2023).
@@
The paper is organized as follows. In Sec.IIII, we introduce the notation and conventions we are going to use. Then, in Sec. $III$, we start with the metric-affine analogue of the theory introduced in Ref. Gottlober et al. (1990), which was studied in the Riemannian setting. Immediately the differences with the metric formalism emerge. By solving the connection field equations of the theory we cast it in an equivalent Scalar-Tensor metric theory with higher order derivative terms for the scalar field. Consequently we generalize it in Sec. $IV$ and study the metric-affine version of the general class of theories with the Lagrangians $\mathcal{L}=RF(\Box)R$ and $\mathcal{L}=RF(\Box^{-1})R$. Deriving the field equations, and solving the connection equations, we again establish equivalence with a generalized higher-derivative Scalar-Tensor theory over a Riemannian background. Finally, we introduce a new non-local Metric-Affine Gravity theory which has no Riemannian analogue and consists of the newly introduced terms $X\Box X$, where $X$ stands for torsion and non-metricity, into the action, along with the usual Einstein-Hilbert term. Obtaining the field equations we show how torsion and non-metricity are dynamical fields in this setting even in vacuum. Some applications to cosmology are discussed in Sec $VI$. In Sec. $VII$, we then wrap up our results and point to future directions.
@@ II The Metric-Affine formalism
The difference between the general connection $\Gamma^{\lambda}{}_{\mu\nu}$ and the Levi-Civita one, $\tilde{\Gamma}^{\lambda}{}_{\mu\nu}$, defines a tensor that quantifies how much the geometry deviates from being Riemannian. This is the so-called distortion tensor Schouten (1954); Hehl et al. (1995a)
@@
This is called the post-Riemannian expansion of the connection and it is particularly useful for splitting the various quantities into their Riemannian parts [^1] Namely the ones computed with respect to the Levi-Civita connection. plus the non-Riemannian contributions of torsion and non-metricity. Torsion and non-metricity can then be recovered from the distortion from the relations (see e.g. Hehl et al. (1995b)),
@@
In this form, it can further act upon tensor fields of arbitrary rank, $\Box T^{\mu_{1}...\mu_{p}}{}{}_{\nu_{1}...\nu_{q}}=\widetilde{\nabla}^{\mu}\widetilde{\nabla}_{\mu}T^{\mu_{1}...\mu_{p}}{}{}_{\nu_{1}...\nu_{q}}$. Note that in Briscese and Pucheu (2016) the definition
@@
where $\nabla$ is the full covariant derivative associated with the full connection $\Gamma^{\lambda}{}_{\mu\nu}$, was used in a Palatini formulation of non-local gravity. Although this could be a viable option, there are several reasons to avoid such a definition in the current context. First and foremost, (10) does not reduce to (8) when acted upon functions. Secondly, given the fact that the metric is not covariantly constant with respect to the full connection, the above definition is not the only one that reduces to $\partial^{\mu}\partial_{\mu}$ in the flat space limit. Indeed, operators like $\nabla_{\mu}(g^{\mu\nu}\nabla_{\nu})$, $\frac{1}{\sqrt{-g}}\nabla_{\mu}(\sqrt{-g}g^{\mu\nu}\nabla_{\nu})$ or $\nabla_{\mu}(\nabla_{\nu}g^{\mu\nu})$ also reduce to $\partial^{\mu}\partial_{\nu}$ in the flat space limit. One would then need to take all possible linear combinations of such operators (with the coefficients added up to one) in order not to discriminate among them, see for instance Shimada et al. (2019). However, this approach, not only does it complicate the analysis but also introduces new arbitrary parameters [^2] As an example, let us consider the linear combination $\Box=ag^{\mu\nu}\nabla_{\mu}\nabla_{\nu}+b\nabla_{\mu}g^{\mu\nu}\nabla_{\nu}$. Consistency in the flat space limit then demands that $a+b=1$. Furthermore using the connection decomposition (5) it follows that $\Box f=\Box_{g}f+\Big(\frac{Q^{\mu}}{2}+2S^{\mu}+(b-1)q^{\mu}\Big)\partial_{\mu}f$. That is, it can always be split into the Riemannian box plus derivative couplings of the function and the non-Riemannian vectors.. Thirdly, for operators like (10), terms like $R_{\mu\nu}\Box^{(\Gamma)}R^{\mu\nu}$ and $R^{\mu\nu}\Box^{(\Gamma)}R_{\mu\nu}$ do not coincide since the metric cannot freely move through the full covariant derivative $\nabla_{\mu}$. This then complicates the analysis even further since one would be forced to include both of them in order to maintain full generality. Finally, the basic property of the inner product $(\phi\Box\psi)=(\Box\phi,\psi)$ is lost if one uses (10) instead of (8). It is therefore only natural to stick to (8) for the definition of the box operator in our current context. [^3] The same definition of the box operator was also used in Bombacigno et al. (2025).
@@ III A simple Warm-up Theory
Let us start with the model in Gottlober et al. (1990), but now in the metric-affine formalism, where the connection and the metric are independent fields. We have
@@
where $\alpha$ is dimensionless [^4] Of course, the General Relativity limit is recovered for $\alpha=1$. and $\beta$ and $\gamma$ are parameters with mass dimension 2-2 and 4-4 respectively. Also, we define the box operator associated to the metric $g$, in the usual way, (see eq. (8)) as we explained in the previous section. Let us also note that, in this action, the scalar curvature $R$ is now the generic one that has also torsion and non-metricity. The action (11) is, up to surface terms, equivalent to [^5] Of course when acting all scalars, the covariant derivative, of any connection, always reduces to the partial one Schouten (1954). Therefore, for the scalar curvature we have $\nabla_{\mu}R=\widetilde{\nabla}_{\mu}R=\partial_{\mu}R$.
@@
Using the method developed in Iosifidis (2019), from the above connection field equations, one finds the distortion
@@
and consequently, torsion and non-metricity read
$S_{\mu\nu}{}{}^{\lambda}=\delta^{\lambda}_{[\mu}\left(\frac{1}{2}q_{\nu]}-\frac{1}{(n-2)}\frac{\partial_{\nu]}\phi}{\phi}\right)$ (18)
$Q_{\nu\alpha\mu}=\frac{1}{2}g_{\alpha\mu}q_{\nu}$ (19)
Let us observe that both are of purely vectorial type. Now notice that our action (11) is invariant under projective transformations of the connection
@@
since it is built solely from $R$ which is itself projectively invariant. This means that there is an undetermined vectorial degree of freedom which can be fixed by a gauge choice Iosifidis (2019) [^6] It should be noted that this invariance has implications when matter is added. In particular, it forces the dilation part of the hypermomentum to be vanishing (see for instance Hehl et al. (1981)). Of course this constraint is avoided if the matter does not couple to the connection but only to the metric.. Two special gauge choices are of interest here. The first one is to fix
@@
This is the vanishing torsion gauge. We see therefore a torsion/non-metricity duality similar to the one that appears in $f(R)$ theories Iosifidis et al. (2019). Let us now investigate the metric field equations (13). Taking the trace of the latter, it follows that
@@
with $\phi=1+2\gamma\Box\psi$. We see therefore that the initial theory is on-shell equivalent to a Scalar-Tensor theory with higher order derivatives for the scalar field. It is important to stress that, in contrast to the Riemannian case where the resulting metric field equations contain sixth order derivatives of the scalar field, in the metric-affine approach developed here, we have up to fourth-order derivatives of the scalar field. In order words, with respect to the metric formalism developed in Gottlober et al. (1990) [^7] See also Amendola et al. (1993) for a cosmological application., the metric-affine approach reduces the order of the derivatives describing the dynamics of the extra scalar field.
@@ IV.1 Nonlocal Ultraviolet $\mathcal{L}=R+RF(\Box)R$ Theories
where $F(\Box)$ is an analytic function of the box operator, usually called the form factor Bas i Beneito et al. (2024), which we may expand as
@@
In the first representation $M^{2}$ carries a squared mass scale and consequently the coefficients $a_{m}$ are dimensionless constants. In the second form, the mass dependence is absorbed in $f_{m}$. In order to have more simplified expressions, we shall use the latter representation. Non-local theories of this kind, namely consisting of positive powers of the box operator are relevant in the UV regime Buoninfante et al. (2022). In the metric formalism, it was shown in Biswas et al. (2006) that the class (50) gives rise to ghost and asymptotically free theories. Below we shall analyze the dynamics of these theories in the metric-affine approach.
@@
Note that these are formally the same with (14), with the mere replacement $\gamma\Box R\rightarrow F(\Box)R$ and of course setting $\alpha=1$ and $\beta=0$. It is therefore again possible to solve for the connection. Indeed, despite the complexity of the metric field equations, the connection field equations obtained above, have a rather simple structure. As a matter of fact, since the form of (47) is identical to that of (14) with only the exact form of $\phi$ altered, the solution to (47) is still
@@ IV.2 Non-local Infrared $\mathcal{L}=R+RF(\Box^{-1})R$ Theories
The previous results can be straightforwardly applied to non-local theories involving the inverse box operator Deser and Woodard (2007) $\Box^{-1}$ defined through $\Box\Box^{-1}=\Box^{-1}\Box=\mathbb{1}$. As a prototype let us consider the class of non-local theories given by [^8] For the study of such theories in the metric formalism see e.g. Capozziello et al. (2020); Capozziello and Capriolo (2021); Capozziello et al. (2023).
@@
where $\Box^{-m}$ is an abbreviation for $(\Box^{-1})^{m}$. Note that this form of series for $F$, namely containing powers of the inverse box operator are relevant for the IR regime. In order to proceed with the field equations, we only need to vary the defining relation of $\Box^{-1}$ and subsequently multiply from the left by the same operator, to trivially get [^9] We assume that a Green’s function and boundary conditions are fixed and no homogeneous modes of $\Box$ appear.
@@ V Novel Theories with no Riemannian analogue
As we have already pointed out, the action (11) as well as its extension (50), have been well studied in the Riemannian case. In particular, in a Riemannian setting, (11) is the most general theory that is quadratic in scalar curvature. We note, however, that in MAG, squared torsion and non-metricity terms supplement the linear in curvature Einstein-Hilbert term. Therefore, in a MAG scheme before extending to (11) it would be quite natural to consider non-local actions consisting of scalars formed by sandwiching the box operator with torsion and non-metricity before moving to quadratic curvature. Given that there are 11 parity even quadratic invariants (see Pagani and Percacci (2015) and also Iosifidis and Koivisto (2019) for the proof that this is the number of independent invariants), we get the following scalars
Pure Non-Metricity Scalars
$A_{1}=Q_{\alpha\mu\nu}Q^{\alpha\mu\nu}$ (56)
$A_{2}=Q_{\alpha\mu\nu}Q^{\mu\nu\alpha}$ (57)
$A_{3}=Q_{\mu}Q^{\mu}$ (58)
$A_{4}=q_{\mu}q^{\mu}$ (59)
$A_{5}=Q_{\mu}q^{\mu}$ (60)
Pure Torsion Scalars
$B_{1}=S_{\alpha\mu\nu}S^{\alpha\mu\nu}$ (61)
$B_{2}=S_{\alpha\mu\nu}S^{\mu\nu\alpha}$ (62)
$B_{3}=S_{\mu}S^{\mu}$ (63)
Mixed
$C_{1}=Q_{\alpha\mu\nu}S^{\alpha\mu\nu}$ (64)
$C_{2}=Q_{\mu}S^{\mu}$ (65)
$C_{3}=q_{\mu}S^{\mu}\,$, (66)
and then the analogue of the theory (11) to MAG is
@@
Some comments are in order. The first two lines consist of the usual Einstein-Hilbert term along with the quadratic invariants in torsion and non-metricity. These exhaust the number of parity even invariants up to dimension $L^{-2}]$. When only the first two lines of the above action are considered then the vacuum theory reduces to General Relativity Iosifidis (2022a). The third and fourth lines are the new non-local terms we consider. These have not been studied previously in the literature. Let qualitatively study their dynamics. The full field equations are quite lengthy and are given in the appendix. Suppressing indices, the connection field equations schematically look like
@@
where $N$ is the distortion. This shows that torsion and non-metricity have dynamics in this case and the theory has more degrees of freedom than General Relativity. It goes beyond the scope of the current study to investigate the healthy sub-sectors of this class of theories. However, it is worth mentioning that the non-local terms of the third and fourth lines, when added to the quadratic curvature MAG theories, may help to extend the healthy spectrum of the quadratic MAG. In order to get our point across, let us study a simple, but yet non-trivial, subsector of the above class.